A strip of silicone of area 800nm carries a current of 190 Ma calculate current density
1
Expert's answer
2020-10-27T10:12:59-0400

Solution

Current density is given by

"\rho =\frac {I}{A}"


Putting the value of current and area

"\rho =\frac {190\times 10^6}{800\times10^{-9}}"


"\rho=2.375\times10^{14} A/m^2"